# Prediction of Energy Resolution in the JUNO Experiment

### v1

@@ 4 Full detector simulation
The JUNO detector simulation software has been developed based on GEANT4 [34, 35, 36] version 10.04.p02. The simulation software, along with other offline data processing modules, is implemented within the SNiPER (Software for Non-collider Physics Experiment) [37, 38] framework. Further details are outlined in [39]. Aspects that potentially influence the energy resolution which are modeled by the detector simulation are discussed in the following sections. Aspects include: the detector geometry, particle interactions in the detector media, light production, LS optical model describing the light propagation, and PMT optical model accounting for the reflection and angular-dependent PDE of the PMTs.
@@ 4.1 Detector geometry modelling
In detector simulation, mechanical design drawings, survey data and the best-known values for the optical properties of materials and component dimensions are at the basis to enable reliable simulation of light propagation, collection, and the detector non-uniformity response. The key components modelled include:
@@
  • Acrylic sphere with an inner radius of 17.7 m and a thickness of 12.4 cm.
@@
  • Detailed LPMT geometry [17], including photocathode, reflective aluminum film, inner electrode, and supporting structures, which are important for tracking photons inside the LPMTs. The 17,612 LPMTs are positioned in the inner water buffer facing the acrylic sphere, with a minimum distance of 1.42 m between acrylic and the center of the photocathode. In addition, the acrylic and SS protection covers [23] are also modeled, located at the front and the back of the LPMTs, respectively.
  • 25,600 SPMTs [40] located within the gaps between LPMTs with a position distribution arranged to guarantee a uniform SPMT density.
@@
  • The PMT modules which optically isolate the inner and outer water pools..
@@
The optical properties of each key component in the CD have been well defined in the simulation. In Figure 3a, the RI of the acrylic, water and PMTs’ glass (Pyrex) is summarized as a function of wavelength. The RI ($n$) of acrylic and glass is measured at 5 different wavelengths, indicated as markers. These measured values are then extrapolated to other wavelengths using the dispersion relation:
@@
where $\lambda$ is the wavelength. $p_{0}$ and $p_{1}$ are two parameters obtained by fitting the 5 measured data points using Eq. 5. RI below 300 nm is not set in the simulation, because a negligible number of photons from LS are expected to reach the acrylic or PMTs in this wavelength region, due to the strong absorption of LS. The RI of water is obtained from measurements in [41]. The attenuation length of acrylic is obtained by analyzing the transmittance data published in [42, 43], shown as the red curve in Fig. 3b. The attenuation length of water, shown as the blue curve in Fig. 3b, is assumed to be 40 m at 430 nm and its dependence with wavelength is provided by the Daya Bay collaboration. The absorption length is not set for the PMTs’ glass, since this effect is implicitly included in the PMTs’ PDE. The reflectivity of the SS components is assumed to be 53.5% without angular and spectral dependence, which is calculated using Fresnel’s equations with the refractive index obtained from [44]. This quantity will be measured in situ at JUNO and will be updated in the simulation in the future. The optical properties and modelling of the LS and PMTs will be discussed in more detail in Sec. 4.4 and 4.6, respectively.
Figure 3: (a) Refractive index of acrylic (red), PMT glass (green), and water (blue) as a function of wavelength. (b) Attenuation length of acrylic (red) and water (blue) as a function of wavelength. Markers indicate measurements, while solid lines are interpolations or extrapolations.
@@ 4.2.1 Energy deposition
The low energy Livermore model [45] is chosen to simulate the electromagnetic processes of electrons, positrons, gammas, hadrons, and ions in the JUNO simulation. These processes include the photo-electric effect, Compton scattering, Rayleigh scattering, gamma conversion, Bremsstrahlung, and ionisation. The Livermore model is expected to possess improved simulation accuracy of energy deposition and secondary particle generation compared to the standard electromagnetic models, because it directly uses shell cross-section data, instead of parameterizations of these data. The Livermore model can handle particle interactions with energies down to 10 eV and is valid for elements with atomic number between 1 and 99. The physics list construction of G4EmLivermorePhysics provided by GEANT4 is directly used in the JUNO simulation. The production cuts are set to 1 mm for gammas and 0.1 mm for electrons and positrons, which are used to determine the energy thresholds of secondary particle generation during GEANT4 tracking. These cuts can significantly influence the determination of the Birks’ coefficient. Beside the low energy electromagnetic processes, a complete physics list has been constructed in simulation, including hadronic processes, ion physics, lepton and gamma-nuclear interactions, absorption, short-lived particles, radioactive decays, and optical processes.
@@ 4.2.2 Determination of Birks’ coefficient
Constraining the Birks’ coefficient ($kB$) through the energy non-linearity response of an LS-based detector, such as those measured by the Daya Bay and Borexino detectors [46, 47], can be challenging. This is because $kB$ is strongly correlated with the Cherenkov contribution, as discussed in the Daya Bay publication [46]. However, it is possible to determine $kB$ by fitting the LS non-linearity data obtained from table-top measurements. These measurements typically involve electrons with energies below the Cherenkov threshold, typically around 0.2 MeV. Nevertheless, the $kB$ values reported in publications are not directly applicable for use in detector simulations, because when extracting $kB$ using Eq. 1, only the $dE/dx$ of the primary particle is considered, usually calculated using tools like ESTAR [48] or SRIM [49]. This calculation does not account for the production of secondary particles. However, in the JUNO detector simulation, a significant number of secondary particles can be generated, carrying a fraction of the primary particle’s energy. These secondary particles are independently tracked in GEANT4. To address this issue, a new fitting method has been explored to determine $kB$, which can be directly used in simulation. This fitting method takes into account the production and distribution of energy among secondary particles during simulation.
We start by simulating electron interactions in the LS with energies up to 0.2 MeV. The production cuts in simulation are kept the same with those used in JUNO. It is important to note that the choice of production cuts can affect the energy deposition and, consequently, the value of $kB$. Therefore, it is crucial to keep the production cuts consistent throughout the study to ensure the reliability of the results. During the simulation, the energy deposition in each step is recorded. The quenching energy, $E_{q}$, for each event is calculated by summing up all the steps in the event as indicated by Eq. 6, where $S$ is a fitting parameter and represents a normalization factor. A combined fit is then performed on the collected experimental data on electron quenching by minimizing the value of the $\chi^{2}$ function Eq. 7:
$E_{q}$ $=\sum\limits_{\textrm{step}}\frac{S\cdot dE}{1+kB\left(\frac{dE}{dx}\right)}$ (6)
$\chi^{2}$ $=\sum_{n}\sum_{i}\left(\frac{\bar{E}_{q}-M_{i}^{n}}{\sigma_{i}^{n}}\right)^{2},$ (7)
@@
where $n$ denotes the number of data sets, $i$ refers to the number of data points in each data set, $\bar{E}_{q}$ is the average quenched energy of the simulation samples at the corresponding electron energy, and $M$ indicates the value of the experimental data with its corresponding error denoted by $\sigma$. Without loss of generality, we scale the $E_{\textrm{vis}}/E_{\textrm{dep}}$ ratio to 1 at 0.1 MeV for each data set. Figure 4 shows the combined fit result on the electron quenching effect. In these data sets, the electrons come from the Compton scattering of the incident gammas emitted by radioactive sources. The gamma sources are placed outside a cylindrical LS container and can be rotated around it. The data set marked with the green color is from [50], in which a ${}^{22}\textrm{Na}$ source is used and only energies below 0.2 MeV are considered. The data set shown in blue is taken by the JUNO collaborators of Technical University of Munich (TUM), in which the gamma source of ${}^{137}\textrm{Cs}$ is rotated to change its position, and four different data sets are obtained, corresponding to four different rotation angles. Fig. 4 shows one data set while the remaining data sets are used to estimate the systematic errors. From the fitting, Birks’ parameter $kB$ is found to be $(12.1\pm 0.3)\times 10^{-3}\textrm{g}/\textrm{cm}^{2}/\textrm{MeV}$.
@@
Figure 4: Combined fitting results on the electron quenching data below 0.2 MeV with production cuts of 1 mm for gammas and 0.1 mm for electrons and positrons. The data set shown in brown color is from [50]. The data set in blue is taken by the JUNO collaborators of TUM, in which the systematic errors are estimated by measuring the quenching curves at four different incident angles of a ${}^{137}\textrm{Cs}$ gamma source with respect to the LS sample. The red curve is the result of the fit using Eq. (7).
@@ 4.3 Cherenkov radiation
The RI of LS are crucial parameters for calculating the number of Cherenkov photons using Eq. 2. In our simulation, we use the LS RI data shown in Fig. 5, which were obtained with a precision better than 0.01% at five different wavelengths (indicated as markers) using the V-prism refractometer [51]. Then, we employ the dispersion relation given by Eq. 5 to extend the RI values down to 200 nm. For the wavelength range between 120 nm and 200 nm, we adopt the RI shape from the KamLAND experiment [52] and scale it to match the RI of the JUNO LS at 200 nm. Finally, we perform a linear extrapolation to estimate the RI values down to 80 nm, which has a value close to unity.
Figure 5: RI of LS as a function of wavelength. The markers represent the data points measured by experiments using the V-prism refractometer [51]. Then, the dispersion relation (Eq. 5) is used to extend the RI down to 200 nm (yellow region). For wavelength range between 120 nm and 200 nm (gray region), the RI shape is from the KamLAND experiment. For wavelengths below 120 nm (green region), a linear extrapolation is utilized.
Cherenkov photon production is handled by GEANT4, although with some modifications. This is necessary because GEANT4 assumes a RI monotonically increasing with photon energy, where the maximum RI corresponds to the maximum photon energy. However, this is not the case for the LS RI. To address this, we have enhanced the Cherenkov process in GEANT4 to handle more general forms of RI vs photon energy curves. Initially, we use the photon energy range of the RI that are above the Cherenkov threshold based on the velocity information of incident particles. Subsequently, we calculate the number of Cherenkov photons for each energy range using Eq. 2. Finally, we sample the energies of the emitted Cherenkov photons according to the LS RI curve.
It is essential to acknowledge that the RI and re-emission probability of the LS carry significant uncertainties, especially at shorter wavelengths such as the vacuum ultraviolet region ($<200$ nm). These uncertainties introduce a potential bias in predicting the Cherenkov light yield. To address this, we introduce a Cherenkov light yield factor, denoted as $f_{C}$, which is used to adjust the Cherenkov light yield in the simulation. The Cherenkov light yield factor is applied as follows:
@@
Here, $N_{\textrm{CherenOP}}^{\textrm{G4}}$ represents the calculated Cherenkov photons number obtained from GEANT4 using the LS RI as input. The factor $f_{C}$ is determined by constraining it with the LS energy non-linearity and energy scale measurements performed by the Daya Bay detectors. Further details regarding this constraint will be discussed in Sec. 4.5.
@@ 4.4 LS optical model and optical properties
Photon propagation in the LS is governed by the LS optical model, which takes into account the processes of emission, scattering, absorption, and re-emission. In this model, the three components of the LS, namely LAB, PPO, and bis-MSB, are treated as a collective entity and share a set of equivalent optical parameters. These parameters include the photon emission spectrum, absorption and scattering lengths, and the quantum yield after photon absorption. The LS optical model is illustrated in Fig. 6. During the propagation of photons, they can either be absorbed or scattered by the LS, depending on the absorption and scattering lengths and their respective energy. If a photon is absorbed without undergoing the re-emission process, its trajectory is terminated. However, if re-emission occurs, a new photon is generated, and its energy is sampled from the LS emission spectrum. This newly generated photon continues its propagation within the LS.
The reflection and transmittance at the interfaces between the LS and the acrylic, as well as between the acrylic and the inner water buffer, or between the water and the PMT glass bulb, are accounted for by GEANT4 using the Fresnel formula and the predefined RI of the LS, acrylic, water, and glass (Fig. 3a). Additionally, photons have a probability of being absorbed by the acrylic or the water, determined by their respective absorption lengths. Scattering within water uses Rayleigh scattering lengths calculated from the RI by GEANT4. Scattering within acrylic is currently neglected due to the lack of scattering length information. The boundary processes at the surfaces of the other detector components are also described by GEANT4, utilizing predefined optical properties such as reflectivity. After photon propagation, a fraction of the generated photons can impinge upon the PMT photocathodes, where they are further handled by the PMT optical model discussed in Sec. 4.6.
Figure 6: The schematic diagram of light propagation in the LS optical model.
@@
  • Emission spectrum: the LS optical model employs the emission spectrum of bis-MSB, considering the large size of the JUNO detector. After undergoing several cycles of absorption and re-emission processes, the scintillation photons are expected to shift towards longer wavelengths and be primarily dominated by the emissions from bis-MSB, rather than PPO fluor. The bis-MSB emission spectrum is measured using a Fluorolog Tau-3 spectrometer, as shown in Fig.2 of [53]. This spectrum is used to sample the energies of the scintillation photons during their production induced by ionization excitation and the re-emission process.
@@
  • Absorption length: the attenuation length ($L_{att}$) of the JUNO LS is assumed to have a target value of 20 m at 430 nm, which comprises both the absorption length ($L_{abs}$) and the scattering length ($L_{sca}$). This can be expressed as $\frac{1}{L_{att}(\lambda)}=\frac{1}{L_{sca}(\lambda)}+\frac{1}{L_{abs}(\lambda)}$. Subtracting the measured scattering length from the attenuation length yields an absorption length of 77 m at 430 nm. The wavelength dependence of the absorption length is assumed to be the same as that of the Daya Bay LS [57].
  • Quantum yield: the spectrum of the quantum yield is assumed from the Daya Bay experiment as shown in Fig. 7, which has been fine-tuned to achieve agreement between the LS energy non-linearity in data and the simulation. The LS replacement experiment at Daya Bay [21] indicates that the differences in PPO and bis-MSB concentrations between the Daya Bay LS (3 g/L PPO and 15 mg/L bis-MSB) and the JUNO LS have a negligible impact on the quantum yield.

### v2

@@ 4 Full detector simulation
The JUNO detector simulation software was developed based on GEANT4 [41, 42, 43] version 10.04.p02. The simulation software, along with other offline data processing modules, is implemented within the Software for Non-collider Physics Experiment (SNiPER) [44, 45] framework. Further details are outlined in [46]. Aspects that potentially influence the energy resolution, which are modeled by the detector simulation, are discussed in the following sections. The aspects include the detector geometry, particle interactions in the detector media, light production, LS optical model describing the light propagation, and PMT optical model accounting for the reflection and angular-dependent PDE of the PMTs.
@@ 4.1 Detector geometry modelling
In detector simulation, mechanical design drawings, survey data, and the best-known values for the optical properties of materials and component dimensions are crucial to enable reliable simulation of light propagation, collection, and the detector non-uniformity response. The key modelled components include the following:
@@
  • Acrylic sphere with an inner radius of 17.7 m and thickness of 12.4 cm.
@@
  • Detailed LPMT geometry [21], including photocathode, reflective aluminum film, inner electrode, and supporting structures, which are important for tracking photons inside the LPMTs. The 17,612 LPMTs are positioned in the inner water buffer facing the acrylic sphere, with a minimum distance of 1.42 m between acrylic and the center of the photocathode. The acrylic and SS protection covers [29] are also modeled, located at the front and back of the LPMTs, respectively.
  • 25,600 SPMTs [26] located within the gaps between LPMTs with a position distribution that guarantees a uniform SPMT density.
@@
  • PMT modules that optically isolate the inner and outer water pools..
@@
The optical properties of each key component in the CD have been well defined in the simulation. In Figure 3a, the refractive indices of the acrylic, water and PMTs’ glass (Pyrex) are summarized as a function of wavelength. The refractive index of acrylic and glass is measured at 5 different wavelengths, indicated as markers. These measured values are then extrapolated to other wavelengths using the dispersion relation:
@@
where $\lambda$ is the wavelength. $p_{0}$ and $p_{1}$ are two parameters obtained by fitting the 5 measured data points using Eq. 5. Refractive index below 300 nm is not set in the simulation because a negligible number of photons from LS are expected to reach the acrylic or PMTs in this wavelength region, due to the strong absorption of LS. The refractive index of water was obtained from measurements in [47]. The attenuation length of acrylic was obtained by analyzing the transmittance data published in [48, 49], shown as the red curve in Fig. 3b. The attenuation length of water, shown as the blue curve in Fig. 3b, is assumed to be 40 m at 430 nm, and its dependence on wavelength was taken from the Daya Bay collaboration. The absorption length is not set for the PMT glass because this effect is implicitly included in the PMTs’ PDE. The reflectivity of the SS components is assumed to be 53.5% without angular or spectral dependence, which is calculated using Fresnel’s equations with the refractive index obtained from [50]. This quantity will be measured in situ at JUNO and will be updated in the simulation in the future. The optical properties and modelling of the LS and PMTs will be discussed in more detail in Sec. 4.4 and 4.6, respectively.
Figure 3: (a) Refractive indices of acrylic (red), PMT glass (green), and water (blue) as a function of wavelength. (b) Attenuation lengths of acrylic (red) and water (blue) as a function of wavelength. Markers indicate measurements, while solid lines are interpolations or extrapolations.
@@ 4.2.1 Energy deposition
The low energy Livermore model [51] is chosen to simulate the electromagnetic processes of electrons, positrons, gammas, hadrons, and ions in the JUNO simulation. These processes include the photo-electric effect, Compton scattering, Rayleigh scattering, gamma conversion, Bremsstrahlung, and ionisation. The Livermore model is expected to possess improved simulation accuracy of energy deposition and secondary particle generation compared to the standard electromagnetic models because it directly uses shell cross-section data instead of parameterizations of these data. The Livermore model can handle particle interactions with energies down to 10 eV and is valid for elements with atomic number between 1 and 99. The physics list construction of G4EmLivermorePhysics provided by GEANT4 is directly used in the JUNO simulation. The production cuts are set to 1 mm for gammas and 0.1 mm for electrons and positrons, which are used to determine the energy thresholds of secondary particle generation during GEANT4 tracking. These cuts can significantly influence the determination of the Birks’ coefficient. Beside the low-energy electromagnetic processes, a complete physics list has been constructed in simulation, including hadronic processes, ion physics, lepton and gamma-nuclear interactions, absorption, short-lived particles, radioactive decays, and optical processes.
@@ 4.2.2 Determination of Birks’ coefficient
Constraining the Birks’ coefficient ($kB$) through the energy non-linearity response of an LS-based detector, such as those measured by the Daya Bay and Borexino detectors [52, 53], can be challenging. This is because $kB$ is strongly correlated with the Cherenkov contribution, as discussed in the Daya Bay publication [52]. However, it is possible to determine $kB$ by fitting the LS non-linearity data obtained from table-top measurements. These measurements typically involve electrons with energies below the Cherenkov threshold, typically around 0.2 MeV. Nevertheless, the $kB$ values reported in publications are not directly applicable to detector simulations because when extracting $kB$ using Eq. 1, only the $dE/dx$ of the primary particle is considered, usually calculated using tools like ESTAR [54] or SRIM [55]. This calculation does not account for the production of secondary particles. However, in the JUNO detector simulation, a significant number of secondary particles can be generated, carrying a fraction of the primary particle’s energy. These secondary particles are independently tracked in GEANT4. To address this issue, a new fitting method has been explored to determine $kB$, which can be directly used in simulation. This fitting method takes into account the production and distribution of energy among secondary particles during simulation.
@@
The electron quenching effect of JUNO LS has been investigated by two groups: the Institute of High Energy Physics (IHEP) and the Technical University of Munich (TUM). Both groups used electrons generated from the Compton scattering of incident gammas emitted by radioactive sources. The gamma sources are placed outside a cylindrical LS container and can be rotated around it. The data-set marked with brown color was measured by the IHEP group [56], in which a ${}^{22}\textrm{Na}$ source was used and only energies below 0.2 MeV were considered. The data set shown in blue was collected by the TUM group, where the gamma source of ${}^{137}\textrm{Cs}$ was rotated to change its position, and four different data-sets were obtained, corresponding to four different rotation angles. In Fig. 4, only one TUM data-set is shown, while the remaining data-sets are used to estimate the systematic errors.
At each data point in Fig. 4, electron interactions in the LS are simulated using GEANT4, with the initial electron energy $E_{\textrm{true}}$ derived from the Compton-scattered gamma. The production cuts in simulation are kept the same as those used in JUNO. It is important to note that the choice of production cuts can affect the energy deposition and consequently the value of $kB$. Therefore, it is crucial to keep the production cuts consistent throughout the study to ensure the reliability of the results. During the simulation, the energy deposition in each step is recorded. The visible energy, $E_{\textrm{vis}}$, for each event is calculated by summing up all the steps in the event, as indicated by Eq. 6, where $S$ is a fitting parameter and represents a normalization factor. A combined fit is then performed on the collected experimental data on electron quenching by minimizing the value of the $\chi^{2}$ function in Eq. 7:
$E_{\textrm{vis}}$ $=\sum\limits_{\textrm{step}}\frac{S\cdot dE}{1+kB\left(\frac{dE}{dx}\right)}$ (6)
$\chi^{2}$ $=\sum_{n}\sum_{i}\left(\frac{\bar{E}_{\textrm{vis}}/E_{\textrm{true}}-M_{i}^{n}}{\sigma_{i}^{n}}\right)^{2},$ (7)
where $n$ denotes the number of data-sets, $i$ refers to the number of data points in each data-set, ${E}_{\textrm{vis}}$ is the average visible energy of the simulation samples at the corresponding electron energy $E_{\textrm{true}}$, and $M$ indicates the measured ratio of $E_{\textrm{vis}}$ to $E_{\textrm{true}}$, with its corresponding error denoted by $\sigma$. Without loss of generality, we scale the $E_{\textrm{vis}}/E_{\textrm{true}}$ ratio to 1 at 0.1 MeV for each data-set. From the fitting, the Birks’ parameter $kB$ is determined to be $(12.1\pm 0.3)\times 10^{-3}\textrm{g}/\textrm{cm}^{2}/\textrm{MeV}$, with the uncertainty including both statistical and systematic errors.
@@
Figure 4: Combined fitting results on the electron quenching data below 0.2 MeV with production cuts of 1 mm for gammas and 0.1 mm for electrons and positrons. The data-set shown in brown color is from [56]. The data-set in blue is from the JUNO collaborators of TUM, where the systematic errors were estimated by measuring the quenching curves at four different incident angles of a ${}^{137}\textrm{Cs}$ gamma source with respect to the LS sample. The red curve is the fitting result using Eq. (7).
@@ 4.3 Cherenkov radiation
The refractive index of LS is crucial for calculating the number of Cherenkov photons using Eq. 2. In our simulation, we use the LS refractive index data shown in Fig. 5, which were obtained with a precision better than 0.01% at five different wavelengths (indicated as markers) using the V-prism refractometer [57]. Then, we employ the dispersion relation given by Eq. 5 to extend the refractive index down to 200 nm. For the wavelength range between 120 nm and 200 nm, we adopt the refractive index shape from the KamLAND experiment [58] and scale it to match the refractive index of the JUNO LS at 200 nm. Finally, we perform a linear extrapolation to estimate the refractive index down to 80 nm, which has a value close to unity.
Figure 5: Refractive index of LS as a function of wavelength. The markers represent the data points measured by experiments using the V-prism refractometer [57]. Then, the dispersion relation (Eq. 5) is used to extend the refractive index down to 200 nm (yellow region). For wavelength range between 120 nm and 200 nm (gray region), the refractive index shape is taken from the KamLAND experiment. For wavelengths below 120 nm (green region), a linear extrapolation is utilized.
Cherenkov photon production is handled by GEANT4 but with some modifications. This is necessary because GEANT4 assumes a refractive index that monotonically increases with photon energy, where the maximum refractive index corresponds to the maximum photon energy. However, this is not the case for the LS refractive index. To address this, we have enhanced the Cherenkov process in GEANT4 to handle more general forms of refractive index vs. photon energy curves. Initially, we use the photon energy range of the refractive index above the Cherenkov threshold based on the velocity information of incident particles. Subsequently, we calculate the number of Cherenkov photons for each energy range using Eq. 2. Finally, we sample the energies of the emitted Cherenkov photons according to the LS refractive index curve.
It is essential to acknowledge that the refractive index and re-emission probability of the LS carry significant uncertainties, especially at shorter wavelengths, such as in the vacuum ultraviolet region ($<200$ nm). These uncertainties introduce a potential bias in predicting the Cherenkov light yield. To address this, we introduce a Cherenkov light yield factor, denoted as $f_{C}$, which is used to adjust the Cherenkov light yield in the simulation. The Cherenkov light yield factor is applied as follows:
@@
Here, $N_{\textrm{CherenOP}}^{\textrm{G4}}$ represents the calculated Cherenkov photon number obtained from GEANT4 using the LS refractive index as input. The factor $f_{C}$ is determined by constraining it with the LS energy non-linearity and energy scale measurements performed by the Daya Bay detectors. Further details regarding this constraint are discussed in Sec. 4.5.
@@ 4.4 LS optical model and optical properties
Photon propagation in the LS is governed by the LS optical model, which takes into account the processes of emission, scattering, absorption, and re-emission. In this model, the three components of the LS, namely, LAB, PPO, and bis-MSB, are treated as a collective entity and share a set of equivalent optical parameters. These parameters include the photon emission spectrum, absorption and scattering lengths, and quantum yield after photon absorption. The LS optical model is illustrated in Fig. 6. During the propagation of photons, they can either be absorbed or scattered by the LS, depending on the absorption and scattering lengths and their respective energy. If a photon is absorbed without undergoing the re-emission process, its trajectory is terminated. However, if re-emission occurs, a new photon is generated, and its energy is sampled from the LS emission spectrum. This newly generated photon continues its propagation within the LS.
The reflection and transmittance at the interfaces between the LS and acrylic, as well as between the acrylic and inner water buffer or between the water and PMT glass bulb, are accounted for by GEANT4 using the Fresnel formula and the predefined refractive indices of the LS, acrylic, water, and glass (Fig. 3a). Additionally, photons have a probability of being absorbed by the acrylic or water, determined by their respective absorption lengths. Scattering within water uses Rayleigh scattering lengths calculated from the refractive index by GEANT4. Scattering within acrylic is currently neglected due to the lack of scattering length information. The boundary processes at the surfaces of the other detector components are also described by GEANT4, utilizing predefined optical properties such as reflectivity. After photon propagation, a fraction of the generated photons can impinge upon the PMT photocathodes, where they are further handled by the PMT optical model discussed in Sec. 4.6.
Figure 6: Schematic diagram of light propagation in the LS optical model.
@@
  • Emission spectrum: Considering the large size of the JUNO detector, the LS optical model employs the emission spectrum of bis-MSB. After undergoing several cycles of absorption and re-emission processes, the scintillation photons are expected to shift towards longer wavelengths and be primarily dominated by the emissions from bis-MSB, rather than PPO fluor. The bis-MSB emission spectrum is measured using a Fluorolog Tau-3 spectrometer, as shown in Fig.2 of [59]. This spectrum is used to sample the energies of the scintillation photons during their production induced by ionization excitation and the re-emission process.
@@
  • Absorption length: The attenuation length ($L_{att}$) of the JUNO LS is assumed to have a target value of 20 m at 430 nm, which comprises both the absorption length ($L_{abs}$) and scattering length ($L_{sca}$). This can be expressed as $\frac{1}{L_{att}(\lambda)}=\frac{1}{L_{sca}(\lambda)}+\frac{1}{L_{abs}(\lambda)}$. Subtracting the measured scattering length from the attenuation length yields an absorption length of 77 m at 430 nm. The wavelength dependence of the absorption length is assumed to be the same as that of the Daya Bay LS [63].
  • Quantum yield: The spectrum of the quantum yield was taken from the Daya Bay experiment, as shown in Fig. 7, which has been fine-tuned to achieve agreement between the LS energy non-linearity in data and the simulation. The LS replacement experiment at Daya Bay [25] indicates that the differences in PPO and bis-MSB concentrations between the Daya Bay LS (3 g/L PPO and 15 mg/L bis-MSB) and JUNO LS have a negligible impact on the quantum yield.